OpenStudy (anonymous):

Question 3.3. Jeff invests an amount at 4% interest compounded annually. After 3 years, he has $1687.30. What was the original amount Jeff invested? $750 $1200 $1450 $1500 Question 4.4. If $3000 is deposited in an account that pays 5% interest, what is the difference in the amount after 4 years between the amount earned if the principal is compounded annually and the amount earned calculated using simple interest? $30.72 $41.12 $46.52 $53.76

OpenStudy (anonymous):

a = p(1 + r/100)^3 1687.30 = p(1+ 4/100)^3